Find the length of a rectangular lot with a perimeter of 136 meters if the length is 6 meters more than the width
Answer by ikleyn(52864)

.
37 m and 31 m.

You can solve it mentally:

L + W = 68     (= 136/2)
L - W =  6.
-------------Add the two equations . . .